a crystal is made of particles x,y,z.x forms fcc packing,y occupies all the octahedral voids of x and z occupies all the tetrahedral voids of x.the formula of the crystal is
1]x2yz2
2]xyz2
3]x8y4z5
4]x5y4z8

Grade:8

2 Answers

Shivam jha
6 years ago
For X in fcc-6×1/2=3For Y in octahedral void in fcc arrangment-4For X in tetrahedral voids in fcc arrangment=8 So formula=X3 Y4 Z8
Juri
6 years ago
In a fcc along the one body diagonal, there is one octahedral void and two tetrahedral void.. So initially there was y=4 and z=2x4=8 but after the particles along one body diagonal removed then y is left with 3 and Z is left with 6. Now in x we have 6x1/2+ (8-2)/8 atoms, i.e. x=3+3/4=15/4 Therefore, compound will be x5y4z8